# Towards Bidirectional Human-AI Alignment: A Systematic Review for Clarifications, Framework, and Future Directions

### v2

@@ 3.2. Scope
Alignment. Our review encompassed all the alignment goals outlined in Section 3.1.2. Since many studies emphasized the importance of value alignment (Gabriel, 2020; Russell, 2014), we particularly summarized a clarified taxonomy of alignment values and identified the value in each paper (if applicable). Additionally, we focus on analyzing the AI models’ output and generation, but not the neural network’s intermediate representations (McGrath et al., 2023; Park et al., 2023; Xie et al., 2024), for alignment research.
3.3. Systematic Literature Review Process
3.3.1. Identification and Screening with Keywords.
We started with papers published in the AI-related domain venues (including NLP, HCI, and ML fields) beginning from the advent of general-purpose generative AI to present, i.e., primarily between January, 2019 and January, 2024 (see details in Appendix B.1). We retrieved 34,213 papers in the initial Identification stage. Further, we collectively defined a list of keywords (see details in Appendix B.2) and screened for papers that included at least one of these keywords (e.g., human, alignment) or their variations either in the title or abstract. We included 2,136 papers in this Screening stage.
3.3.2. Assessing Eligibility with Criteria
We further filtered the 2,136 papers based on explicit inclusion and exclusion criteria, i.e., the Eligibility stage. Our criteria revolved around six research questions that we collectively identified to be most pertinent to the topic, including 1) what essential human values have been aligned by some AI models? 2) how did we effectively quantify or model human values to guide AI development? 3) what strategies have been employed to integrate human values into the AI development process? 4) how did existing studies improve human understanding and evaluation of AI alignment? 5) what are the practices for designing interfaces and interactions that facilitate human-AI collaboration? 6) How have AI been adapted to meet the needs of various human value groups? We included papers that could potentially answer any of these questions. Further, based on the scope in Section 3.2, we excluded papers that did not meet our inclusion criteria. This resulted in a final corpus of 411 papers, which were analyzed in detail using qualitative coding (see Appendix B.3 for more details).
3.3.3. Qualitative Code Development.
Referring to the code development process in Lee et al. (2024), we first conducted qualitative coding for each paper by identifying relevant sentences that could answer the above research questions, and entering short codes to describe them into a codebook. We iteratively coded relevant sentences from each paper through a mix of inductive and deductive approaches, which allowed flexibility to expand, modify or change the driving research questions based on our learnings as we went through the process. To ensure rigor in our coding process, two authors coded each paper. The first author independently annotated all papers after reviewing the paper abstracts and introductions. Twelve team members each annotated a subset of the paper corpus. Our corpus includes papers from different domains (e.g., HCI, NLP and ML). Therefore, we divided the authors into HCI and NLP/ML333Note that NLP and ML are two different domains, we combine them together for the purposes of literature review analysis since they both work on developing and evaluating AI technologies. teams and assigned the papers accordingly based on expertise. All team members coded each of their assigned papers to answer all six questions (if applicable) introduced above.
3.3.4. Framework Development and Rigorous Coding.
After developing annotations, all authors collaborated to create the bidirectional human-AI alignment framework by integrating the annotations within each of the codes. The initial version of the framework was proposed by the author who reviewed all papers. This framework furthermore underwent iterative improvement through: 1) discussions with all team members involved in paper coding, and 2) revisions based on feedback from the project advisors. Additionally, we strengthened the framework by reviewing papers from the AI Ethics conferences (including FAccT and AIES), and related work of the collected papers that covered other domains such as psychology and social science. We further added missing codes and papers to ensure comprehensive coverage (see Appendix B.1 for details). The final bidirectional human-AI alignment framework, with detailed topologies, is presented in Section 4. Following the framework’s finalization, we conducted another separate coding process to annotate whether each paper investigated dimensions within our framework. Two authors independently coded each paper.444The joint probability of agreement for the paper annotations was 0.78. These codes were then used to perform quantitative and qualitative analyses, as presented in Section 5.
@@ 4. Bidirectional Human-AI Alignment Framework
This section introduces the Bidirectional Human-AI Alignment framework developed from the systematic literature review.
To encompass the ongoing, mutual process of human-AI alignment, we design the framework to include: “Align AI to Humans” and “Align Humans to AI”.
As shown in Figure 1, research in the
“Align AI to Humans” direction studies mechanisms to ensure that AI systems’ values match those of humans’ (Section 4.1 and 4.2). In comparison, studies in the
“Align Humans to AI” direction investigate the humans’ cognitive and behavioral adaptation to the AI advancement (Section 4.3 and 4.4).
To probe into the research challenges and existing exploration of the solutions in literature, we further developed the fine-grained typologies (in Figure 6 and Figure 7) to extend the high-level framework in Figure 1.
@@
ALIGN AI to HUMANS
Studies in this direction aim to integrate human specifications to train, steer and customize AI systems. The two main challenges involved in this direction include: carefully specifying the values of the system, and ensuring that system adopts the specification robustly (Wikipedia, 2024; Ngo et al., 2024). Therefore, as shown in Figure 6, we design the two core research questions in this direction as: RQ1. Human Values and Specifications (Section 4.1) and RQ2. Integrating Human Specifications into AI (Section 4.2)
4.1. Align AI to Humans:
@@
To clarify and systematically understand human values relevant to human-AI alignment, we leverage the adapted “Schwartz Theory of Basic Values” introduced in Section 3.1.3, and examine the category of aligned human values from the two dimensions of “Sources” and “Types”.
@@
Sources of Values. This dimension examines the three sources of human values (Schwartz, 1994). Individual sources indicate values from individuals comprise universal needs of individuals as biological organisms (Schwartz, 1994) or prioritize personal interests (Kiesel et al., 2022). At this level, value alignment can be usually assessed independently of the context of interaction. This perspective highlights the values about technical capabilities of AI models, including factuality (Roit et al., 2023; Hu et al., 2024a), calibration (Liu et al., 2024a), output diversity (Bradley et al., 2024), and model inductive bias (Si et al., 2023a). Additionally, this code includes research on aligning model behaviors with the characteristics and preferences of individual humans, e.g., predict human moral judgements and decisions (Jin et al., 2022), and cognitive biases (Jones and Steinhardt, 2022; Koo et al., 2023). Social sources mean values from the social groups include universal requirements for smooth functioning and survival of groups (Schwartz, 1994). Value alignment at this level transcends personal interactions and emphasizes the broader categories defined by shared experiences, identities, cultures, norms, and more (Santy et al., 2023). Research in this area often targets the alignment of AI behaviors with the general preferences of humans (Hendrycks et al., 2021; Bakker et al., 2022; Lee et al., 2023a). Additionally, there are efforts aimed to align AI with specific targeted groups, examining issues through the lenses of fairness (Fleisig et al., 2023), social norms (Sky et al., 2023), morality (Prabhumoye et al., 2021; Rao et al., 2023), and beyond. Interactive sources considers values at the interaction level to include universal requisites of coordinated social interaction (Schwartz, 1994, 2012), which typically occurs in interpersonal situations, such as in the dynamics of speaker-receiver relationships during language communication (Hovy and Yang, 2021). In the context of “human-AI alignment”, we adjust the definition of Interaction values to be the AI values that humans expect for AI as tools, such as Usability (Zhang et al., 2023a), Autonomy (Wang et al., 2023a), among others. Research along this line focuses on alignment strategies to enhance human-AI interaction (Gupta et al., 2023; Oh et al., 2023), collaborative decision-making (Gajos and Mamykina, 2022; Yang et al., 2023a), and trust (Buçinca et al., 2021; Dhuliawala et al., 2023).
@@
Types of Values. In this dimension, we introduce the five high-order human values derived from a combination of the Schwartz Human Values (Schwartz, 1994) and empirical values studies from the surveyed papers. We provide a more in-depth taxonomy of value types in Section 5.2, including the relationship of these five types and more nuanced value categories. Self-Enhancement refers to a set of self-protective and personal values that emphasize enhancing self-esteem and a sense of personal worth (Sedikides and Strube, 1995). One of the most important aspect is achievement, is competence as judged by social standards, which includes general capability (effectiveness/efficiency) as covered by the majority of the research (Guo et al., 2024; Chen et al., 2023b; Yan et al., 2023a). Another dimension is power, which relates to social status and prestige, as well as control or dominance over people and resources (Ngo et al., 2024; Pinski et al., 2023). Self-Transcendence refers to a set of self-expanding and socially-focused human values that emphasize expanding beyond oneself (Frankl, 1966). One critical aspect is benevolence, which relates to preservation and enhancement of the welfare of people with whom one is in frequent personal contact. Extensive efforts have investigated the values including helpfulness (Askell et al., 2021; Bai et al., 2022a), honesty/factuality (Roit et al., 2023; Hu et al., 2024a), responsibility/accountability (Smith-Renner et al., 2020a; Solyst et al., 2023a). Another critical aspect is universalism, which relates to understanding, appreciation, tolerance and protection for the welfare of all people and for nature. Researchers have looked into values including inclusion/broad-mindedness (Bakker et al., 2022; Dong et al., 2024) and equality/fairness (Shen et al., 2022b; Wang et al., 2023b). Conservation refers to a set of self-protective and socially-focused human values that hold and safeguard traditional institutions and customs (Hamilton, 2020). Under conservative values, people are concerned about security (Ssafety, harmony, and stability of society, relationships, and oneself), tradition (Respect for and acceptance of the customs and ideas provided by traditional culture or religion), and conformity (restraint of actions, inclinations, and impulses likely to upset or harm others and violate social expectations or norms). Exemplar studies in this field include safety (Qi et al., 2023; Dai et al., 2024; Anwar et al., 2024), mental health (Liu et al., 2023a), and cultural moral norms (Ramezani and Xu, 2023). Openness to Change refers to a set of self-expanding and personally-focused human values motivated by an anxiety-free need to grow, in contrast to conservation. Under this type, people are concerned about stimulation (seeking excitement, novelty, and challenges in life), hedonism (pursuing pleasure and sensuous gratification for oneself), and self-direction (independent thought and action—choices, creativity, and exploration). Exemplar studies in this field include understanding of privacy (Mireshghallah et al., 2024; Yao et al., 2024) and creativity (Isajanyan et al., 2024; Ashkinaze et al., 2024). We adopt the four high-order human values from Schwartz Human Values (Schwartz, 1994), supplementing them with additional values empirically collected from survey papers as shown in Figure 4. Additionally, as the conventional theory is missing the context of AI-system and human-AI interactions, we introduce a new high-order value type called Desired Values for AI Tools. This category encompasses the values humans expect from AI when used as tools in “human-AI interaction”. These values include usability (competence according to the human experience on AI functionality) and human-likeness (resemblance to human intelligence and behavior). Exemplary studies in this area include assessments of AI usability (Zhang et al., 2023a) and autonomy (Wang et al., 2023a).
@@
This sub-research question investigates how human values are interactively555 We define ”interactions” broadly to encompass both ”synchronous” and ”asynchronous” interactions between humans and AI: (1) Synchronous Interactions: These are real-time exchanges where humans and AI systems interact simultaneously. Examples include live chatbots, virtual assistants, and real-time decision-making systems where immediate responses are required. (2) Asynchronous Interactions: These interactions do not occur in real-time, allowing for delays between actions and responses, such as data annotations by humans, and any AI system that processes human inputs and provides outputs after a certain period. By including both synchronous and asynchronous interactions, we aim to cover the full spectrum of ways in which humans and AI systems can communicate and collaborate. specified for AI systems to ensure alignment. It aims to elucidate the interaction techniques by which AI systems manifest or instantiate human values, thereby revealing the underlying mechanisms that shape their behavior or functionality.
@@
Explicit Human Feedback. This dimension refers to the direct specification of human values through explicitly defined formats or mechanisms. Principles provides AI systems with explicitly defined principles, guidelines, or rules that dictate behavior or decision-making in alignment with human values (Petridis et al., 2024a). Rating and Ranking is widely used to assign numerical scores or rankings to options or outcomes based on their alignment with human values (Dai et al., 2024). Natural Language Interaction/Conversations allows humans to interact with AI systems through natural language interfaces to express and communicate human values (Bai et al., 2022b). For Multimodal Feedback, human values can also be provided in multiple modalities, such as sketches/images and gestures (Chung et al., 2022), to convey human values.
@@
Implicit Human Feedback. This dimension refers to the indirect representation or inference of human values within AI systems through patterns, signals, or cues embedded in the data or decision-making processes. Discarded Options refers to the options or choices that human discard when interacting with AI systems during the decision-making processes, which also potentially infer human values. (Park et al., 2021b). Language Analysis means that textual data and language patterns can also contain rich information to identify implicit references to human values or value-related concepts (Lu et al., 2023a). Theory of Mind refers to the ability of agents and people to attribute mental states, such as beliefs, intentions, desires, emotions, knowledge, percepts, and non-literal communication, to themselves and others (Ma et al., 2023c). Social Relationships refers to the implicit values derived from analyzing human social relations and behaviors, which are derived from external sources (e.g., social network) that can inherently reflect their values (Isajanyan et al., 2024).
@@
Simulated Human Value Feedback. When human values in explicit formats are expensive or impossible to collect, one may simulate human-like feedback within AI systems to approximate human responses and preferences regarding specific values. Human Feedback Simulators uses computational algorithms to simulate human-like feedback on values, based on predefined criteria or training data (Petridis et al., 2024b; Ouyang et al., 2022; Li et al., 2023c). Comparison to Human Data refers to developing techniques that assess the likelihood or probability of AI-generated outputs matching human behaviors in a reference set or dataset (Dong et al., 2024). Synthetic Data curates data by generating synthetic comparisons based on naive assumptions or heuristic rules, followed by post-validation to ensure feedback quality (Kim et al., 2023a).
4.2. Align AI to Humans:

### v3

@@ 3.4. Scopes and Key Components in Alignment
Alignment. Our review encompassed all the alignment goals outlined in Section 3.1. Since many studies emphasized the importance of value alignment (Gabriel, 2020; Russell, 2014), we particularly summarized a clarified taxonomy of alignment values and identified the value in each paper (if applicable). Additionally, we focus on analyzing the AI models’ output and generation, but not the neural network’s intermediate representations (McGrath et al., 2023; Park et al., 2023; Xie et al., 2024), for alignment research.
3.5. Systematic Literature Review and Framework Rationale
Achieving the alignment between AI and human values necessitates collaboration among interdisciplinary AI practitioners (e.g., developers and researchers). To comprehensively analyze the research literature pertaining to this ongoing, mutual process of human-AI alignment, we conducted a systematic literature review adhering to the the PRISMA guideline (Page et al., 2021; Stefanidi et al., 2023). Our iterative selection process yielded 411 papers from an initial pool of 34,213 publications, spanning the period from January 2019 to January 2024. These papers were sourced from leading venues across multiple AI-related domains, including Natural Language Processing, Human-Computer Interaction, Machine Learning, and more, using well-defined keywords and criteria. A group of 13 interdisciplinary researchers coded the papers. Through iterative discussions, we developed the paper codes and proposed framework (Section 3). Detailed steps and processes are elucidated in Appendix A.
3.6. Rationale Behind the Framework
Our systematic review revealed that current research largely overlooks the nature of long-term dynamic changes and the importance of interaction in human-AI alignment. First, the manifestation of certain AI risks, harms, and the practical quality of AI services primarily emerge through interaction between humans and AI systems (Weidinger et al., 2023). For instance, Shen and Huang (2020) discovered that AI explanations, deemed effective by AI developers and researchers, were found not useful for humans in real-world scenarios of interactive AI error detection. Additionally, the use of AI influences human thinking and behavior, which in turn alters their values when interacting with AI. For instance, Li et al. (2024b) found that incorporating AI assistance in writing workflows changed humans’ writing perceptions, behavior, and performance, creating new requirements for developing the next generation of AI writing assistants (Lee et al., 2024).
In light of these findings, we posit that AI practitioners must expand their focus beyond merely “Aligning AI to Humans” – a paradigm centered on integrating human specifications to train, steer, and customize AI. Equal emphasis must be placed on “Aligning Humans to AI”, a complementary paradigm investigating human cognitive and behavioral adaptation to AI. This latter aspect is crucial in supporting individuals and society at large in understanding, critiquing, collaborating with, and adapting to transformative AI advancements. To encapsulate these comprehensive perspectives, we propose the “Bidirectional Human-AI Alignment” framework, and provide its definition below Figure 1. This holistic framework acknowledges the reciprocal nature of alignment, addressing both the advancements of AI systems to incorporate human values, and the evolving human adaptation of AI systems in real-world practice.
@@ 4. Bidirectional Human-AI Alignment Framework
This section introduces the Bidirectional Human-AI Alignment framework, which encompassing two interconnected alignment directions, as shown in Figure 1.
The
“Align AI to
Humans” direction studies mechanisms to ensure that AI systems’ values match those of humans’ (Section 4.1 and 4.2).
We identify two crucial research questions to be addressed in this direction: RQ1. What relevant human values are studied for AI alignment, and how do humans specify these values? and RQ2. How can human values be integrated into the development of AI?
In comparison, the
“Align Humans to AI” direction investigates the humans’ cognitive and behavioral adaptation to the AI advancement (Section 4.3 and 4.4). We recognize two imperative research questions along this direction: RQ3. How might humans learn to perceive, explain, and critique AI? and RQ4. How do humans and society make behavioral changes and react to AI advancement?.
We address the four questions by elucidating the fine-grained typologies and literature solutions derived from our systematic review, as delineated in Figure 2 and Figure 3, respectively.
@@
ALIGN AI to
HUMANS
This direction delineates alignment research from AI-centered perspective (e.g., ML/NLP domains) and provides AI developers and researchers with approaches for two main challenges: carefully specifying the values of the system, and ensuring that system adopts the specification robustly (Wikipedia, 2024; Ngo et al., 2024). Therefore, as shown in Figure 2, we explore the two core research questions in this direction as: RQ1. Human Values and Specifications (Section 4.1) and RQ2. Integrating Human Specifications into AI (Section 4.2)
4.1. Align AI to
Humans:
@@ 4.1. Align AI to
To clarify and systematically understand human values relevant to human-AI alignment, we leverage the adapted “Schwartz Theory of Basic Values” introduced in Section 3.3, and examine the category of aligned human values from the two dimensions of “Sources” and “Types”.
@@
Sources of Values. This dimension examines the three sources of human values (Schwartz, 1994). Individual sources indicate values from individuals comprise universal needs of individuals as biological organisms (Schwartz, 1994) or prioritize personal interests (Kiesel et al., 2022; Jin et al., 2022). At this level, value alignment can be usually assessed independently of the context of interaction. This perspective highlights the values about technical capabilities of AI models, including factuality (Roit et al., 2023; Hu et al., 2024), calibration (Liu et al., 2024a), output diversity (Bradley et al., 2024), and model inductive bias (Si et al., 2023). Additionally, this code includes research on aligning model behaviors with the characteristics and preferences of individual humans, e.g., predict human moral judgements and decisions (Jin et al., 2022), and cognitive biases (Jones and Steinhardt, 2022; Koo et al., 2023). Social sources mean values from the social groups include universal requirements for smooth functioning and survival of groups (Schwartz, 1994; Ammanabrolu et al., 2022). Value alignment at this level transcends personal interactions and emphasizes the broader categories defined by shared experiences, identities, cultures, norms, and more (Santy et al., 2023). Research in this area often targets the alignment of AI behaviors with the general preferences of humans (Hendrycks et al., 2021; Bakker et al., 2022; Lee et al., 2023a). Additionally, there are efforts aimed to align AI with specific targeted groups, examining issues through the lenses of fairness (Fleisig et al., 2023), social norms (Sky et al., 2023), morality (Prabhumoye et al., 2021; Rao et al., 2023), and beyond. Interactive sources considers values at the interaction level to include universal requisites of coordinated social interaction (Schwartz, 1994, 2012), which typically occurs in interpersonal situations, such as in the dynamics of speaker-receiver relationships during language communication (Hovy and Yang, 2021; Cabrera et al., 2023). In the context of “human-AI alignment”, we adjust the definition of Interaction values to be the AI values that humans expect for AI as tools, such as Usability (Zhang et al., 2023a), Autonomy (Wang et al., 2023a), among others. Research along this line focuses on alignment strategies to enhance human-AI interaction (Gupta et al., 2023; Oh et al., 2023), collaborative decision-making (Gajos and Mamykina, 2022; Yang et al., 2023a), and trust (Buçinca et al., 2021; Dhuliawala et al., 2023).
@@
Types of Values. In this dimension, we introduce the five high-order human values derived from a combination of the Schwartz Human Values (Schwartz, 1994) and empirical values studies from the surveyed papers. We provide a more in-depth taxonomy of value types in Section 5.2, including the relationship of these five types and more nuanced value categories. Self-Enhancement refers to a set of self-protective and personal values that emphasize enhancing self-esteem and a sense of personal worth (Sedikides and Strube, 1995; He et al., 2023; Xu et al., 2024; Huang et al., 2023; Park et al., 2023; Kang et al., 2023; Ye et al., 2024; Lima et al., 2021; Santurkar et al., 2023). One of the most important aspect is achievement, is competence as judged by social standards, which includes general capability (effectiveness/efficiency) as covered by the majority of the research (Guo et al., 2024). Another dimension is power, which relates to social status and prestige, as well as control or dominance over people and resources (Pinski et al., 2023). Self-Transcendence refers to a set of self-expanding and socially-focused human values that emphasize expanding beyond oneself (Frankl, 1966; Banovic et al., 2023; Wu et al., 2023d; Jin et al., 2022; Wu et al., 2023b; Huang et al., 2023; Ma et al., 2023b; Isajanyan et al., 2024; He et al., 2023; Nie et al., 2023; Santurkar et al., 2023; Ramezani and Xu, 2023; Sheng et al., 2021; Goldfarb-Tarrant et al., 2023). One critical aspect is benevolence, which relates to preservation and enhancement of the welfare of people with whom one is in frequent personal contact. Extensive efforts have investigated the values including helpfulness (Bai et al., 2022a), honesty/factuality (Hu et al., 2024), responsibility/accountability (Solyst et al., 2023). Another critical aspect is universalism, which relates to understanding, appreciation, tolerance and protection for the welfare of all people and for nature. Researchers have looked into values including inclusion/broad-mindedness (Bakker et al., 2022; Dong et al., 2024) and equality/fairness (Shen et al., 2022; Wang et al., 2023b). Conservation refers to a set of self-protective and socially-focused human values that hold and safeguard traditional institutions and customs (Hamilton, 2020; Qi et al., 2023; Zagalsky et al., 2021; Lai et al., 2022; Santy et al., 2023; Zhu et al., 2023; Zhou et al., 2023a). Under conservative values, people are concerned about security (Ssafety, harmony, and stability of society, relationships, and oneself), tradition (Respect for and acceptance of the customs and ideas provided by traditional culture or religion), and conformity (restraint of actions, inclinations, and impulses likely to upset or harm others and violate social expectations or norms). Exemplar studies in this field include safety (Anwar et al., 2024), mental health (Liu et al., 2023), and cultural moral norms (Ramezani and Xu, 2023). Openness to Change refers to a set of self-expanding and personally-focused human values motivated by an anxiety-free need to grow, in contrast to conservation (Bakker et al., 2022; Lima et al., 2020; Ashktorab et al., 2020; Jiang et al., 2023; Lima et al., 2020; Kang et al., 2023; Varanasi and Goyal, 2023; Yan et al., 2023; Hosking et al., 2024; Wu et al., 2023d; Kapania et al., 2023). Under this type, people are concerned about stimulation (seeking excitement, novelty, and challenges in life), hedonism (pursuing pleasure and sensuous gratification for oneself), and self-direction (independent thought and action—choices, creativity, and exploration). Exemplar studies in this field include understanding of privacy (Mireshghallah et al., 2024) and creativity (Isajanyan et al., 2024; Ashkinaze et al., 2024). We adopt the four high-order human values from Schwartz Human Values (Schwartz, 1994), supplementing them with additional values empirically collected from survey papers as shown in Figure 2. Additionally, as the conventional theory is missing the context of AI-system and human-AI interactions, we introduce a new high-order value type called Desired Values for AI Tools. This category encompasses the values humans expect from AI when used as tools in “human-AI interaction” (Yang et al., 2023b; Zhang et al., 2023a; Oh et al., 2023; Bakker et al., 2022; Boyd, 2021; Wang et al., 2023a; Paleja et al., 2021). These values include usability (competence according to the human experience on AI functionality) and human-likeness (resemblance to human intelligence and behavior). Exemplary studies in this area include assessments of AI usability (Zhang et al., 2023a) and autonomy (Wang et al., 2023a).
@@
This sub-research question investigates how human values are interactively111 We define ”interactions” broadly to encompass both ”synchronous” and ”asynchronous” interactions between humans and AI: (1) “Synchronous Interactions” indicates real-time exchanges where humans and AI systems interact simultaneously. (2) “Asynchronous Interactions” allows for delays between actions and responses, such as data annotations by humans to train the AI models. specified for AI systems to ensure alignment. It aims to elucidate the interaction techniques by which AI systems manifest or instantiate human values, thereby revealing the underlying mechanisms that shape their behavior or functionality.
@@
Explicit Human Feedback. This dimension refers to the direct specification of human values through explicitly defined formats or mechanisms. Principles provides AI systems with explicitly defined principles, guidelines, or rules that dictate behavior or decision-making in alignment with human values (Petridis et al., 2024a; Bai et al., 2022b). Rating and Ranking is widely used to assign numerical scores or rankings to options or outcomes based on their alignment with human values (Dai et al., 2024; Yuan et al., 2023). Natural Language Interaction/Conversations allows humans to interact with AI systems through natural language interfaces to express and communicate human values (Bai et al., 2022b; Hwang et al., 2023; Guo et al., 2024; Feng and Boyd-Graber, 2019; Yao et al., 2023a; Lin et al., 2024; Hong et al., 2022; Shen et al., 2023a; Zhang et al., 2021). For Multimodal Feedback, human values can also be provided in multiple modalities, such as sketches/images and gestures (Chung et al., 2022), to convey human values (Chen et al., 2022; Geva et al., 2022; Wu et al., 2022; Jörke et al., 2023).
@@
Implicit Human Feedback. This dimension refers to the indirect representation or inference of human values within AI systems through patterns, signals, or cues embedded in the data or decision-making processes. Discarded Options refers to the options or choices that human discard when interacting with AI systems during the decision-making processes, which also potentially infer human values (Park et al., 2021b). Language Analysis means that textual data and language patterns can also contain rich information to identify implicit references to human values or value-related concepts (Lu et al., 2023). Theory of Mind refers to the ability of agents and people to attribute mental states, such as beliefs, intentions, desires, emotions, knowledge, percepts, and non-literal communication, to themselves and others (Ma et al., 2023b). Social Relationships refers to the implicit values derived from analyzing human social relations and behaviors, which are derived from external sources (e.g., social network) that can inherently reflect their values (Isajanyan et al., 2024).
@@
Simulated Human Value Feedback. When human values in explicit formats are expensive or impossible to collect, one may simulate human-like feedback within AI systems to approximate human responses and preferences regarding specific values. Human Feedback Simulators uses computational algorithms to simulate human-like feedback on values, based on predefined criteria or training data (Petridis et al., 2024b; Li et al., 2023a). Comparison to Human Data refers to developing techniques that assess the likelihood or probability of AI-generated outputs matching human behaviors in a reference set or dataset (Dong et al., 2024). Synthetic Data curates data by generating synthetic comparisons based on naive assumptions or heuristic rules, followed by post-validation to ensure feedback quality (Kim et al., 2023).
4.2. Align AI to
Humans: